Valproic acid may be used in the management of epilepsy. Its chemical structure is shown below, it has a pKa of 4.8;



a) Valproic acid may be given orally where it will pass through the different pH environments of the GI tract. Calculate the percentage ionisation of valproic acid in the duodenum and ileum which typically have pH values of 6 and 7.4 respectively (40% of marks)


b) What mass of anhydrous potassium dihydrogen phosphate and dipotassium hydrogen phosphate would you need to manufacture 500ml of 50mM pH 7.0 phosphate buffer that could be used for an injectable formulation of valproic acid. The pKa of dihydrogen phosphate is 7.2 and the RMM of potassium dihydrogen phosphate and dipotassium hydrogen phosphate are 136 and 174 respectively. (60% of marks)
1
Expert's answer
2021-04-22T05:47:03-0400

HA+A=0.05

A/HA=0.063

HA=0.047M

A=0.003M

This means, you need to put 0.047 moles of KH2PO4 and 0.003 moles of K2HPO4 salts into 1 liters of solution.

for molecular weights: KH2PO4= 39+97=136amu; K2HPO4= 39*2+96=174amu

Thus, 0.047*136=6.392g of KH2PO4 and 0.003*174= 0.522g of K2HPO4 should be added.